OpenStudy (mysticfury):

Complete the equation for the linear function whose graph contains the points (9, 7) and (4, –8). y − ____ = ____ (x - 9) @SfunnSneffanS

OpenStudy (anonymous):

Ok, so here there are two points. Do you know have to find the steepness?

OpenStudy (mysticfury):

the slope you mean? and yes i think so

OpenStudy (anonymous):

Yes, the slope :)

OpenStudy (mysticfury):

y^2 - y^1 -------- x^2 - x^1 -8 - 7 ------ 4 - 9 -15 --- -5 3 or -3 ?

OpenStudy (anonymous):

3 is correct, good! So now you know the slope, what remains is to find the intercept with the y-axis, i.e. the constant term

OpenStudy (anonymous):

Now, as we know the slope and one point which the line should go throughm we can do as in the other task: Set the constant term equal to a number a and solve for that number for one of the points (9,7) or (4,-8).

OpenStudy (mysticfury):

okay but what space does it go in? im not used to this formula they are using

OpenStudy (anonymous):

Do you mean in wether the line goes like this / or like this \ ?

OpenStudy (mysticfury):

no i mean for the last question we did this 2 = 0.9 x -1 + x <(solving for this) i understand why you put the numbers like you did in that one but for this im not sure what the equation would be to solve for it like we did above y - ___ = ___ (x - 9) i dont understand that equation ^^

OpenStudy (anonymous):

Ok, I understand. Actually, if we first find the equation of the line on the form y = 3*x + a, we can then afterwards write it on the form y - ___ = ___ (x - 9), finding the the ___ numbers. This will simply be a matter of adding and subtracting from each side of the equation. And no matter how we choose to write the equation, the slope will always be 3. Ok?

OpenStudy (anonymous):

So actually, we already know this ___(x-9) number. It is 3.

OpenStudy (mysticfury):

okay so we know 3(x - 9) ?

OpenStudy (anonymous):

Yes. Now, we can plug in a point, say (9,7). x = 9, y = 7. Let us call the unknown number ___ in y - ___ for c.

OpenStudy (mysticfury):

okay so 7 - c = 3(9 - 9) ?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

So, you can now find that c = 7. Then we have the result y - 7 = 3*(x-9). To make sure we did no mistake, plug in the two points the line should go through and check if both sides of the equation are equal. If they are, we can be certain that we did it right

OpenStudy (mysticfury):

alright ! thank you so much that was my last question :)

OpenStudy (mysticfury):

i wish i could give you more then one medal for being so Helpful and Awesome !

OpenStudy (anonymous):

Ok, no problem. The last thing I mentioned is something you always can do to check if you have found the right equation for the line:) Thank you, very glad I could help you understand :)